OpenStudy (aroub):

It took Wally t minutes to commute to work in the morning. Due to rush hour traffic, his return trip took an extra 18 minutes. Given that the total time Wally spent in his car for both trips was 54 minutes, and that the distance between his home and office is 27 miles, determine the average speed for each leg. Explain the answer :)

OpenStudy (anonymous):

d=vt t=d/v t1, v1 to work t2,v2 to home t1+18=27/v2 t2=27/v2 27/v2-18+27/v2=54 v2=.75 mi/min *60 min/hr = 45mi/hr At this speed, the time required would be t2=d/v=27/45=3/5 hr or 36 minutes. t1=t2-18=36-18=18 v1=d/t1=27mi/18min=1.5mi/min*60min/1hr=90mi/hr

OpenStudy (anonymous):

i forgot to put in that t2=t1+18

OpenStudy (aroub):

Wait, you have to work with me a bit.. What did you exactly do here: t1+18=27/v2 ?

OpenStudy (anonymous):

That's where I forgot to put t2=t1+18, time 2=time 1+18

OpenStudy (anonymous):

so t1=t2-18

OpenStudy (anonymous):

i then added t1 and t2 to get 54 minutes in terms of v2, the average speed home

OpenStudy (anonymous):

it is kind of messy, lol

OpenStudy (aroub):

Haha, it's okay. I got it! Thank you =DD

hero (hero):

 2t + 18 = 54 2t = 54 - 18 2t = 36 t = 36/2 t = 18